# CSI-Bench: A Large-Scale In-the-Wild Dataset for Multi-task WiFi Sensing

### v1

@@ C.1 Evaluation with Difficulty Tiers
LSTM [17] 97.62 ±0.52 97.62 ±0.52 69.12 ±5.63 68.20 ±5.19 67.12 ±2.96 66.05 ±4.03
MLP [32] 94.84 ±0.85 94.84 ±0.85 70.59 ±9.61 70.19 ±9.91 63.70 ±2.37 63.41 ±2.25
PatchTST [29] 97.13 ±0.72 97.13 ±0.72 61.76 ±7.59 56.31 ±12.22 62.67 ±2.82 61.36 ±4.14
ResNet18 [16] 97.27 ±0.32 97.27 ±0.32 77.94 ±5.63 76.96 ±6.46 68.84 ±3.04 68.08 ±3.58
TimeSformer-1D [8] 96.58 ±0.50 96.59 ±0.49 67.65 ±7.59 64.55 ±11.72 65.75 ±9.29 61.19 ±17.16
Transformer [34] 97.08 ±0.54 97.08 ±0.54 69.12 ±5.63 68.10 ±5.91 65.07 ±6.94 63.89 ±7.40
ViT [11] 97.40 ±0.42 97.40 ±0.42 77.94 ±13.04 77.07 ±14.46 65.75 ±3.71 64.06 ±6.66
@@
LSTM [17] 99.11 ±0.17 99.11 ±0.17 98.61 ±0.13 98.61 ±0.13 98.08 ±0.28 98.08 ±0.28
MLP [32] 98.54 ±0.14 98.54 ±0.14 97.67 ±0.15 97.67 ±0.15 96.46 ±0.13 96.46 ±0.13
PatchTST [29] 99.20 ±0.06 99.20 ±0.06 98.77 ±0.19 98.77 ±0.19 98.49 ±0.22 98.49 ±0.22
ResNet18 [16] 98.94 ±0.17 98.94 ±0.17 98.42 ±0.16 98.42 ±0.16 98.32 ±0.25 98.32 ±0.25
TimeSformer-1D [8] 99.05 ±0.22 99.05 ±0.22 98.29 ±0.31 98.29 ±0.31 98.60 ±0.23 98.60 ±0.23
Transformer [34] 98.23 ±0.24 98.23 ±0.24 97.31 ±0.47 97.31 ±0.47 97.54 ±0.31 97.54 ±0.31
ViT [11] 99.56 ±0.08 99.56 ±0.08 99.41 ±0.08 99.41 ±0.08 99.17 ±0.11 99.17 ±0.11
Table 9 presents breathing detection results, where all models maintain high accuracy and F1-scores ($>$96%) across tiers. ViT performs best, achieving over 99% F1 consistently. LSTM and PatchTST follow closely, especially in the Easy setting. Even in the Hard tier, model performance drops only slightly. ResNet18 and TimeSformer-1D also generalize well, with minimal performance variance. The results suggest that breathing patterns are relatively easier to model and robust to environmental changes.
@@
LSTM [17] 99.72 ±0.32 99.75 ±0.29 100.00 ±0.00 100.00 ±0.00 98.31 ±0.50 98.31 ±0.50
MLP [32] 91.36 ±0.93 92.03 ±0.82 96.11 ±1.31 96.18 ±1.29 80.20 ±1.06 80.03 ±1.19
PatchTST [29] 100.00 ±0.00 100.00 ±0.00 99.90 ±0.19 99.95 ±0.10 99.86 ±0.17 99.86 ±0.18
ResNet18 [16] 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00
TimeSformer-1D [8] 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00
Transformer [34] 99.30 ±0.36 99.40 ±0.24 99.90 ±0.19 99.90 ±0.19 98.95 ±0.66 98.95 ±0.66
ViT [11] 99.79 ±0.42 99.82 ±0.35 99.90 ±0.19 99.90 ±0.19 99.50 ±0.23 99.50 ±0.23
@@
LSTM [17] 96.65 ±0.96 96.99 ±0.78 98.79 ±0.11 98.80 ±0.11 96.94 ±0.94 96.94 ±0.95
MLP [32] 98.21 ±0.28 98.29 ±0.18 99.13 ±0.11 99.13 ±0.11 98.19 ±0.36 98.19 ±0.36
PatchTST [29] 98.01 ±0.69 98.28 ±0.54 98.59 ±0.36 98.59 ±0.36 97.49 ±0.71 97.49 ±0.72
ResNet18 [16] 99.86 ±0.11 99.86 ±0.11 99.73 ±0.05 99.73 ±0.05 99.48 ±0.32 99.48 ±0.32
TimeSformer-1D [8] 96.56 ±0.64 96.92 ±0.56 98.68 ±0.18 98.69 ±0.18 97.32 ±0.31 97.31 ±0.32
Transformer [34] 98.73 ±0.62 98.80 ±0.55 98.63 ±0.17 98.63 ±0.17 98.08 ±0.55 98.08 ±0.55
ViT [11] 98.38 ±0.87 98.41 ±0.81 99.27 ±0.32 99.27 ±0.32 98.10 ±0.45 98.10 ±0.45
Table 10 demonstrates that localization is a highly separable task. Most models—including PatchTST, ResNet18, and TimeSformer-1D—achieve perfect scores in the Easy and Medium tiers and retain near-perfect performance in the Hard tier. ViT, Transformer, and LSTM also show strong results (F1 $>$ 98%). MLP consistently underperforms, particularly in the Hard tier (F1: 80.03%), likely due to limited spatial modeling. Overall, most models handle localization with high reliability. These results indicate that CSI-based localization is a highly separable task, and that most temporal or spatially-aware models can solve it with high reliability.
@@ C.2 Evaluation on OOD Splits
LSTM [17] 60.57 ±2.12 57.04 ±2.32 53.65 ±0.89 46.22 ±0.72 53.33 ±2.11 45.70 ±2.01
MLP [32] 56.33 ±1.23 50.79 ±1.11 52.15 ±0.85 43.45 ±1.40 52.06 ±0.54 42.05 ±0.97
PatchTST [29] 61.61 ±1.81 58.05 ±1.54 56.85 ±0.63 49.55 ±0.47 56.44 ±1.47 49.25 ±1.33
ResNet18 [16] 66.21 ±1.96 63.57 ±1.90 57.98 ±0.87 50.90 ±0.96 59.24 ±1.47 52.07 ±1.53
TimeSformer-1D [8] 60.24 ±1.00 55.70 ±1.20 54.65 ±0.93 46.63 ±0.79 54.95 ±0.84 45.74 ±0.79
Transformer [34] 61.82 ±0.95 57.80 ±0.78 54.92 ±0.98 47.17 ±1.12 54.72 ±0.84 46.67 ±1.00
ViT [11] 66.33 ±1.73 63.65 ±1.69 58.87 ±1.12 51.86 ±1.31 59.00 ±1.36 51.48 ±1.26
@@
LSTM [17] 59.25 ±1.69 59.32 ±1.72
MLP [32] 57.31 ±1.61 57.15 ±1.45
PatchTST [29] 60.45 ±1.07 60.56 ±1.17
ResNet18 [16] 68.07 ±1.93 68.21 ±1.97
TimeSformer-1D [8] 60.84 ±0.81 61.00 ±0.79
Transformer [34] 59.94 ±0.77 59.81 ±0.96
ViT [11] 69.37 ±1.53 69.55 ±1.61
@@
LSTM [17] 24.89 ±2.97 24.29 ±3.02 28.64 ±1.08 26.76 ±1.02 29.20 ±0.55 23.83 ±1.34
MLP [32] 28.73 ±0.89 27.31 ±0.84 25.76 ±0.77 20.86 ±1.41 26.19 ±0.57 17.32 ±0.74
PatchTST [29] 28.13 ±2.17 26.60 ±1.38 26.42 ±1.88 25.35 ±1.63 28.86 ±0.67 23.15 ±0.76
ResNet18 [16] 31.19 ±5.81 27.93 ±4.95 30.67 ±3.06 28.01 ±3.78 32.67 ±1.50 27.64 ±2.53
TimeSformer-1D [8] 27.95 ±2.04 25.85 ±2.67 29.73 ±2.99 27.93 ±3.97 31.19 ±0.87 26.98 ±1.48
Transformer [34] 30.68 ±3.11 28.76 ±3.51 29.67 ±1.63 27.12 ±1.79 30.26 ±1.94 25.97 ±2.26
ViT [11] 32.04 ±1.95 30.11 ±2.12 30.83 ±2.36 28.62 ±2.51 31.54 ±1.66 26.94 ±1.77
Tables 12-14 present the performance of supervised models under three cross-domain generalization settings—Cross-Device, Cross-Environment, and Cross-User—for Human Activity Recognition, Human Identification, and Proximity Recognition, respectively. Across all tasks, ViT consistently achieves the highest performance, with the best F1-scores in most OOD settings.

### v2

@@ C.1 Evaluation with Difficulty Tiers
LSTM [LSTM1997] 97.62 ±0.52 97.62 ±0.52 69.12 ±5.63 68.20 ±5.19 67.12 ±2.96 66.05 ±4.03
MLP [MLP] 94.84 ±0.85 94.84 ±0.85 70.59 ±9.61 70.19 ±9.91 63.70 ±2.37 63.41 ±2.25
PatchTST [PatchTST2023] 97.13 ±0.72 97.13 ±0.72 61.76 ±7.59 56.31 ±12.22 62.67 ±2.82 61.36 ±4.14
ResNet18 [kaiming2016] 97.27 ±0.32 97.27 ±0.32 77.94 ±5.63 76.96 ±6.46 68.84 ±3.04 68.08 ±3.58
TimeSformer-1D [TimeSformer2021] 96.58 ±0.50 96.59 ±0.49 67.65 ±7.59 64.55 ±11.72 65.75 ±9.29 61.19 ±17.16
Transformer [attention2017] 97.08 ±0.54 97.08 ±0.54 69.12 ±5.63 68.10 ±5.91 65.07 ±6.94 63.89 ±7.40
ViT [ViT2021] 97.40 ±0.42 97.40 ±0.42 77.94 ±13.04 77.07 ±14.46 65.75 ±3.71 64.06 ±6.66
@@
LSTM [LSTM1997] 99.11 ±0.17 99.11 ±0.17 98.61 ±0.13 98.61 ±0.13 98.08 ±0.28 98.08 ±0.28
MLP [MLP] 98.54 ±0.14 98.54 ±0.14 97.67 ±0.15 97.67 ±0.15 96.46 ±0.13 96.46 ±0.13
PatchTST [PatchTST2023] 99.20 ±0.06 99.20 ±0.06 98.77 ±0.19 98.77 ±0.19 98.49 ±0.22 98.49 ±0.22
ResNet18 [kaiming2016] 98.94 ±0.17 98.94 ±0.17 98.42 ±0.16 98.42 ±0.16 98.32 ±0.25 98.32 ±0.25
TimeSformer-1D [TimeSformer2021] 99.05 ±0.22 99.05 ±0.22 98.29 ±0.31 98.29 ±0.31 98.60 ±0.23 98.60 ±0.23
Transformer [attention2017] 98.23 ±0.24 98.23 ±0.24 97.31 ±0.47 97.31 ±0.47 97.54 ±0.31 97.54 ±0.31
ViT [ViT2021] 99.56 ±0.08 99.56 ±0.08 99.41 ±0.08 99.41 ±0.08 99.17 ±0.11 99.17 ±0.11
Table 10 presents breathing detection results, where all models maintain high accuracy and F1-scores (>>96%) across tiers. ViT performs best, achieving over 99% F1 consistently. LSTM and PatchTST follow closely, especially in the Easy setting. Even in the Hard tier, model performance drops only slightly. ResNet18 and TimeSformer-1D also generalize well, with minimal performance variance. The results suggest that breathing patterns are relatively easier to model and robust to environmental changes.
@@
LSTM [LSTM1997] 99.72 ±0.32 99.75 ±0.29 100.00 ±0.00 100.00 ±0.00 98.31 ±0.50 98.31 ±0.50
MLP [MLP] 91.36 ±0.93 92.03 ±0.82 96.11 ±1.31 96.18 ±1.29 80.20 ±1.06 80.03 ±1.19
PatchTST [PatchTST2023] 100.00 ±0.00 100.00 ±0.00 99.90 ±0.19 99.95 ±0.10 99.86 ±0.17 99.86 ±0.18
ResNet18 [kaiming2016] 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00
TimeSformer-1D [TimeSformer2021] 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00 100.00 ±0.00
Transformer [attention2017] 99.30 ±0.36 99.40 ±0.24 99.90 ±0.19 99.90 ±0.19 98.95 ±0.66 98.95 ±0.66
ViT [ViT2021] 99.79 ±0.42 99.82 ±0.35 99.90 ±0.19 99.90 ±0.19 99.50 ±0.23 99.50 ±0.23
@@
LSTM [LSTM1997] 96.65 ±0.96 96.99 ±0.78 98.79 ±0.11 98.80 ±0.11 96.94 ±0.94 96.94 ±0.95
MLP [MLP] 98.21 ±0.28 98.29 ±0.18 99.13 ±0.11 99.13 ±0.11 98.19 ±0.36 98.19 ±0.36
PatchTST [PatchTST2023] 98.01 ±0.69 98.28 ±0.54 98.59 ±0.36 98.59 ±0.36 97.49 ±0.71 97.49 ±0.72
ResNet18 [kaiming2016] 99.86 ±0.11 99.86 ±0.11 99.73 ±0.05 99.73 ±0.05 99.48 ±0.32 99.48 ±0.32
TimeSformer-1D [TimeSformer2021] 96.56 ±0.64 96.92 ±0.56 98.68 ±0.18 98.69 ±0.18 97.32 ±0.31 97.31 ±0.32
Transformer [attention2017] 98.73 ±0.62 98.80 ±0.55 98.63 ±0.17 98.63 ±0.17 98.08 ±0.55 98.08 ±0.55
ViT [ViT2021] 98.38 ±0.87 98.41 ±0.81 99.27 ±0.32 99.27 ±0.32 98.10 ±0.45 98.10 ±0.45
Table 11 demonstrates that localization is a highly separable task. Most models—including PatchTST, ResNet18, and TimeSformer-1D—achieve perfect scores in the Easy and Medium tiers and retain near-perfect performance in the Hard tier. ViT, Transformer, and LSTM also show strong results (F1 >> 98%). MLP consistently underperforms, particularly in the Hard tier (F1: 80.03%), likely due to limited spatial modeling. Overall, most models handle localization with high reliability. These results indicate that CSI-based localization is a highly separable task, and that most temporal or spatially-aware models can solve it with high reliability.
@@ C.2 Evaluation on OOD Splits
Tables 13-15 present the performance of supervised models under three cross-domain generalization settings—Cross-Device, Cross-Environment, and Cross-User—for Human Activity Recognition, Human Identification, and Proximity Recognition, respectively. Across all tasks, ViT consistently achieves the highest performance, with the best F1-scores in most OOD settings.
@@
LSTM [LSTM1997] 60.57 ±2.12 57.04 ±2.32 53.65 ±0.89 46.22 ±0.72 53.33 ±2.11 45.70 ±2.01
MLP [MLP] 56.33 ±1.23 50.79 ±1.11 52.15 ±0.85 43.45 ±1.40 52.06 ±0.54 42.05 ±0.97
PatchTST [PatchTST2023] 61.61 ±1.81 58.05 ±1.54 56.85 ±0.63 49.55 ±0.47 56.44 ±1.47 49.25 ±1.33
ResNet18 [kaiming2016] 66.21 ±1.96 63.57 ±1.90 57.98 ±0.87 50.90 ±0.96 59.24 ±1.47 52.07 ±1.53
TimeSformer-1D [TimeSformer2021] 60.24 ±1.00 55.70 ±1.20 54.65 ±0.93 46.63 ±0.79 54.95 ±0.84 45.74 ±0.79
Transformer [attention2017] 61.82 ±0.95 57.80 ±0.78 54.92 ±0.98 47.17 ±1.12 54.72 ±0.84 46.67 ±1.00
ViT [ViT2021] 66.33 ±1.73 63.65 ±1.69 58.87 ±1.12 51.86 ±1.31 59.00 ±1.36 51.48 ±1.26
@@
LSTM [LSTM1997] 59.25 ±1.69 59.32 ±1.72
MLP [MLP] 57.31 ±1.61 57.15 ±1.45
PatchTST [PatchTST2023] 60.45 ±1.07 60.56 ±1.17
ResNet18 [kaiming2016] 68.07 ±1.93 68.21 ±1.97
TimeSformer-1D [TimeSformer2021] 60.84 ±0.81 61.00 ±0.79
Transformer [attention2017] 59.94 ±0.77 59.81 ±0.96
ViT [ViT2021] 69.37 ±1.53 69.55 ±1.61
@@
LSTM [LSTM1997] 24.89 ±2.97 24.29 ±3.02 28.64 ±1.08 26.76 ±1.02 29.20 ±0.55 23.83 ±1.34
MLP [MLP] 28.73 ±0.89 27.31 ±0.84 25.76 ±0.77 20.86 ±1.41 26.19 ±0.57 17.32 ±0.74
PatchTST [PatchTST2023] 28.13 ±2.17 26.60 ±1.38 26.42 ±1.88 25.35 ±1.63 28.86 ±0.67 23.15 ±0.76
ResNet18 [kaiming2016] 31.19 ±5.81 27.93 ±4.95 30.67 ±3.06 28.01 ±3.78 32.67 ±1.50 27.64 ±2.53
TimeSformer-1D [TimeSformer2021] 27.95 ±2.04 25.85 ±2.67 29.73 ±2.99 27.93 ±3.97 31.19 ±0.87 26.98 ±1.48
Transformer [attention2017] 30.68 ±3.11 28.76 ±3.51 29.67 ±1.63 27.12 ±1.79 30.26 ±1.94 25.97 ±2.26
ViT [ViT2021] 32.04 ±1.95 30.11 ±2.12 30.83 ±2.36 28.62 ±2.51 31.54 ±1.66 26.94 ±1.77